OpenStudy (anonymous):

let f(x)=2x^4-4x^2. find all c in the interval (-3, 3) such that f'(c)=0 please show all steps

OpenStudy (mathmate):

f(x)=2x^4-4x^2 f'(x)=8x^3-8x=8x(x^2-1) f'(x)=0 for x=0,x=+/-1 => c=-1,0 or 1